Prove the identity:
1/1-tanx + 1/1+tanx =2/cos^2x

Answer by Alan3354(69443)

Prove the identity:
1/(1-tanx) + 1/(1+tanx) = 2/cos^2x
2/(1 - tan^2) = 2/cos^2
1 - tan^2 = cos^2
Not an identity.